OpenStudy (heart_offire):

A catering company charges $300 plus $40 per guest for a wedding. Sarah and Eric do not want to spend more than $5,000 on catering. Write and solve an inequality in terms of the number of guests, g, that can be invited. A) 300 - 40g ≥ 5000; g ≥ 117 B) 300 - 40g ≤ 5000; g ≤ 118 C) 300 + 40g ≤ 5000; g ≤ 117 D) 300 + 40g ≤ 5000; g ≤ 118

OpenStudy (photon336):

@heart_offire what do you think it is?

OpenStudy (heart_offire):

mmmm between c and d

OpenStudy (photon336):

This is a linear equation and we need to add our initial price of 300 dollars to 40 dollars* the number of guests g so yes I agree with you there \[y = m(x)+b \] \[300$+40(g) = y(g)\]

OpenStudy (photon336):

you can actually just solve this inequality the same way you would solve any other equation. \[300+40g \le 5,000\]

OpenStudy (heart_offire):

ok so i subtract 40 from both sides

OpenStudy (photon336):

you subtract 300 from both sides remember 40 has a variable in it so you want to get that by itself.

OpenStudy (photon336):

Let's pretend that the inequality isn't there: can you solve the rest of this. \[40g = 5,000-300\]

OpenStudy (heart_offire):

40g =4700

OpenStudy (photon336):

so what's g?

OpenStudy (heart_offire):

117.5 = g

OpenStudy (heart_offire):

@mathmale